### v1

# Constrained Entropic Unlearning: A Primal-Dual Framework for Large Language Models

Abstract
Large Language Models (LLMs) deployed in real-world settings increasingly face the need to unlearn sensitive, outdated, or proprietary information. Existing unlearning methods typically formulate forgetting and retention as a regularized trade-off, combining both objectives into a single scalarized loss. This often leads to unstable optimization and degraded performance on retained data, especially under aggressive forgetting. We propose a new formulation of LLM unlearning as a constrained optimization problem: forgetting is enforced via a novel logit-margin flattening loss that explicitly drives the output distribution toward uniformity on a designated forget set, while retention is preserved through a hard constraint on a separate retain set. Compared to entropy-based objectives, our loss is softmax-free, numerically stable, and maintains non-vanishing gradients, enabling more efficient and robust optimization. We solve the constrained problem using a scalable primal-dual algorithm that exposes the trade-off between forgetting and retention through the dynamics of the dual variable. Evaluations on the TOFU and MUSE benchmarks across diverse LLM architectures demonstrate that our approach consistently matches or exceeds state-of-the-art baselines, effectively removing targeted information while preserving downstream utility.
1 Introduction
Large Language Models (LLMs) are now foundational to a wide range of applications, from search engines and coding assistants e.g., [20, 12], to medical diagnostics e.g., [19, 15, 34, 35], scientific research e.g., [1, 31], and education e.g., [26]. Their remarkable performance stems from training on vast, diverse corpora of data. However, this training data often contains sensitive, copyrighted, or ethically problematic content, raising concerns around privacy, misinformation, and regulatory compliance. These concerns have led to a growing demand for machine unlearning, the ability to selectively erase the influence of specific training data or knowledge from a deployed model.
Machine unlearning, initially introduced by Cao and Yang [5], poses a fundamental challenge: how can one remove the impact of a small subset of data without retraining the model from scratch? For LLMs, full retraining is prohibitively expensive, especially as models grow in size. Additionally, LLMs must frequently forget information to comply with regulatory mandates (e.g., the “right to be forgotten”  [43]), avoid generating harmful content [41], prevent the leakage of private data [40], or eliminate reliance on copyrighted materials [9].
This has motivated recent algorithmic efforts to approximate unlearning via lightweight fine-tuning techniques, most notably gradient ascent over the forget set [41, 9, 25]. While such methods can suppress the model’s ability to recall or generate content related to the undesired data, they often do so at a steep cost: they degrade model performance on unrelated, retained data. This degradation is particularly problematic when the forget set is much smaller than the retained dataset, as is typically the case in practice. As a result, recovering the lost utility of the model becomes extremely difficult and costly
Prior work has proposed theoretically principled formulations of unlearning as a regularized optimization problem, seeking model parameters that (i) degrade performance on a forget dataset to remove sensitive or undesirable knowledge, and (ii) preserve performance on a retain dataset to maintain utility. However, this framework encounters three key obstacles.
First, regularized formulations treat unlearning as a soft trade-off between forgetting and retention. In doing so, they often drive the forget loss to vanishingly small values (almost zero e.g.,  [4]) in an effort to eliminate all traces of the undesired data. However, this aggressive forgetting is not only computationally inefficient, but also often counterproductive, leading to a sharp degradation in the utility of the model in unrelated inputs.
Second, at the beginning of unlearning, the initial utility of the model is already high, i.e., the retention loss is small. The regularized formulation attempts to decrease the retention loss at the same time as the forgetting loss, but this is unnecessary: Simply keeping the retention loss at its initial level is enough to maintain the original utility of the model. Trying to reduce the retention loss like the forgetting loss can lead to (i) overfitting and bias toward the retained set, and (ii) difficulty in sufficiently lowering the forgetting loss, since the problem becomes a multi-objective optimization with conflicting gradient directions, and as a result, both forgetting and retention losses decrease poorly.
Third, conventional loss functions such as cross-entropy are poorly behaved when repurposed for forgetting. Cross-entropy is unbounded above; thus, maximizing it results in unstable gradients, optimization divergence, and a lack of interpretability. These behaviors hinder convergence and offer little guidance on when forgetting is sufficient.
Our Contributions. (1) We propose a conceptual shift: unlearning should be cast as a constrained optimization problem, where the goal is to forget designated data while explicitly preserving utility on retained data via a formal constraint. This avoids the need for delicate loss balancing and provides strong performance guarantees. (2) We introduce a novel logit-margin flattening loss that encourages uniform output distributions on the forget set, offering a stable, softmax-free alternative to entropy maximization. This loss is convex, bounded, and provides non-vanishing gradients–making it well-suited for large-scale optimization. (3) We develop a scalable primal-dual algorithm that enforces the retention constraint and captures the trade-off between forgetting and utility through the dynamics of the dual variable. The algorithm supports warm starts and dynamic updates, ensuring efficiency even at LLM scale. (4) Finally, we evaluate our method on the TOFU and MUSE benchmarks using both standard metrics and a new LLM-based judge that assesses behavioral divergence.
000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.7Forget Success (higher is better)Model Utility (higher is better)TARGETRETRAINEDGRAD-DIFFDPONPOSIMNPORMUPDU (OURS)
Figure 1: Comparison of different methods on the TOFU dataset on Llama 3.2 3B: Model Utility vs. forget Success; see Section 4 for explanation of the metrics.
Figure 2: Samples from responses of different methods to same prompt in forgetting and retaining evaluations on the MUSE (Books)
1.1 Related Works
Existing methods for LLM unlearning broadly fall into the following categories:
Retraining-based unlearning approaches involve retraining models from scratch or fine-tuning them on datasets excluding the forget set [2]. Although exact retraining provides the most reliable guarantee of unlearning, it is computationally prohibitive, especially for large-scale LLMs, making it impractical for real-world applications.
Gradient-Ascent-based unlearning techniques commonly apply gradient ascent (GA) to the forget set to suppress undesired model behaviors e.g., [17, 41]. However, these methods can cause gradient explosion, necessitating additional measures, such as gradient clipping or specialized loss functions (e.g., modified cross-entropy); to maintain stability, as in [27] and [38], which employ risk-weighted and regularized variants of gradient ascent. Moreover, they often suffer from catastrophic forgetting [45], markedly degrading model utility because of conflicting optimization objectives.
Optimization-based unlearning methods e.g., [45, 11, 6] update model parameters by attaching explicit penalties to the targeted knowledge. The multi-objective variant  [44] builds on the same framework as [45] but first generates several alternate answers, adding non-trivial overhead before the actual unlearning step. [7] modifies only the teacher logits during self-distillation, leaving the student model vulnerable to attacks. Other works replace the loss itself: for example, [6] adopt an inverted-hinge loss; [42] push the output distribution toward uniformity via KL regularization; and [39] optimize exclusively on the forget set, not considering the retain set. While these designs reduce residual accuracy on the forget set, they commonly over-weight the forget objective, degrading performance on retained data.
Representation-based unlearning operates directly on latent embeddings to remove designated knowledge, e.g., [21].  [22] corrupts hidden states with targeted perturbations, while [24] triggers forgetting through embedding-corrupted prompts. Although such interventions can precisely erase the chosen content, they often distort the surrounding semantic geometry, degrading fluency and factual coherence, and are also hard to scale.
Prompt-based and relabeling unlearning remove information by altering prompts or inverting labels e.g., [30, 9]. Logit-level methods, such as [18], first perform the opposite of unlearning and then apply logit differences for unlearning, which is time consuming. Selective logit adjustment [7], which uses a heuristic for token selection, is likewise unreliable. Although these techniques avoid heavy fine-tuning, they remain brittle: minor paraphrases or adversarial queries can still surface the suppressed knowledge, revealing limited robustness [29].
All existing methods prioritize unlearning performance on the forget set. Overoptimization towards minimizing forget loss can disproportionately degrade model utility, which is challenging to recover given the extensive nature of retain sets. As shown in Figure 2, even retraining the model on the retain data set does not achieve an extremely low loss on the forget set. This suggests that pushing the forget loss lower leads to overforgetting, which makes it easier for later attacks to relearn the forgotten data, since forgetting is not uniform [14]. Another benefit of our method is that it enables a dynamic approach towards handling conflicting gradients. That is, when optimizing two or more losses [41], it is likely that gradients will be conflicting and finding a shared descent direction can be extremely resource intensive for large scale LLMs [27]. However, the primal dual algorithm naturally adjusts the linear combination such that if the retain loss is below our desired threshold, its conflicting dynamic with the forgetting loss could be completely ignored. The dynamicaly changing $\lambda$ would promptly shift attention back to the retention loss if the loss exceeds the user-defined threshold. This approach allows the model to unlearn the forget set while remaining as close as possible to the original model in utility.
1.2 Notation
We denote a general LLM by $\pi$ where $\pi$ takes as input a tokenized prompt $x\in\mathbb{R}^{N\times D}$, where $N=|x|$ is the number of tokens and $D$ is the embedding dimension. The model then outputs $\pi^{\mathrm{logits}}(x)\in\mathbb{R}^{N\times V}$, where $V$ is the vocabulary size of the tokenizer that will be used to decode the tokens into legible text prompts. We let $\pi(x)=\mathrm{Softmax}(\pi^{\mathrm{logits}}(x))$, where the $\mathrm{Softmax}$ operator is applied on each row of its input to turn the logits into probabilities. When the LLM is parametrized by a finite set of parameters $\theta$, we denote it with $\pi_{\theta}$.
2 Problem Formulation
Let $\pi_{\text{ref}}$ be a language model trained or fine-tuned on a dataset $\mathcal{D}$, partitioned into disjoint subsets $\mathcal{D}_{\text{rtn}}$ (data to retain) and $\mathcal{D}_{\text{fgt}}$ (data to unlearn), where each example $(x,y)\in\mathcal{D}$ consists of a prompt $x$ and target response $y$. The objective is to construct a new model $\pi$ that preserves the behavior of $\pi_{\text{ref}}$ on $\mathcal{D}_{\text{rtn}}$ while eliminating knowledge of $\mathcal{D}_{\text{fgt}}$. Perfect unlearning entails eliminating all information related to $\mathcal{D}_{\text{fgt}}$, not merely reducing $\pi(y|x)$, the likelihood of generating $y$ given $x$.
Ideal unlearning would involve retraining a model $\pi_{r}$ from scratch on $\mathcal{D}_{\text{rtn}}$, fully excluding $\mathcal{D}_{\text{fgt}}$. However, this is computationally intensive, costly, and time-consuming, especially given the potential frequency of unlearning requests (e.g., due to outdated data or copyright concerns). Thus, practical unlearning seeks to derive a model $\pi$, close to $\pi_{\text{ref}}$, with the influence of $\mathcal{D}_{\text{fgt}}$ effectively removed.
Unlearning is typically posed as a bi-objective optimization problem that balances the removal of information related to $\mathcal{D}_{\text{fgt}}$ with the preservation of performance on $\mathcal{D}_{\text{rtn}}$. We define two loss functions: $\mathcal{L}_{\text{fgt}}$ to enforce forgetting, and $\mathcal{L}_{\text{rtn}}$ to maintain utility. A common approach is linear scalarization:
$\min_{\pi\in\Pi}\;\mathcal{L}_{\text{fgt}}(\pi,\mathcal{D}_{\text{fgt}})+\lambda\,\mathcal{L}_{\text{rtn}}(\pi,\mathcal{D}_{\text{rtn}}),$ (1)
where $\lambda>0$ controls the trade-off and $\Pi$ denotes a compact function class, e.g., $\{\pi:\|\pi\|_{L_{2}}\leq M\}$. The losses are defined as
$\mathcal{L}_{\text{fgt}}(\pi,\mathcal{D}_{\text{fgt}})=\mathbb{E}_{(x,y)\in\mathcal{D}_{\text{fgt}}}[\ell_{\mathrm{fgt}}(\pi,x,y)],\quad\mathcal{L}_{\text{rtn}}(\pi,\mathcal{D}_{\text{rtn}})=\mathbb{E}_{(x,y)\in\mathcal{D}_{\text{rtn}}}[\ell_{\mathrm{rtn}}(\pi,x,y)],$
where $\ell_{\mathrm{fgt}}$ and $\ell_{\mathrm{rtn}}$ are task-specific loss functions detailed in later sections.
3 Constrained Entropic Unlearning
The linear scalarization formulation in (1) suffers from a limitation: if the forget–set $\mathcal{D}_{\text{fgt}}$ and retain-set $\mathcal{D}_{\text{rtn}}$ overlap (statistically or semantically), blindly reducing $\mathcal{L}_{\text{fgt}}$ can catastrophically increase $\mathcal{L}_{\text{rtn}}$. To balance this trade-off, the scalarization weight $\lambda$ must be carefully tuned for each instance. However, even such a dynamic approach provides no explicit control over the degradation on the retention set. In particular, small values of $\lambda$ may lead to incomplete forgetting, while large values can overly compromise retention.
In light of these challenges, we reformulate unlearning as a constrained optimization problem:
$\begin{split}\min_{\pi\in\Pi}&\quad\mathcal{L}_{\text{fgt}}(\pi,\mathcal{D}_{\text{fgt}})\\ \mathrm{s.t.}&\quad\mathcal{L}_{\text{rtn}}(\pi,\mathcal{D}_{\text{rtn}})\leq\varepsilon.\end{split}$ (2)
Here, $\varepsilon$ is a user-specified threshold for allowable performance degradation on $\mathcal{L}_{\text{rtn}}$. A natural instantiation is
$\varepsilon=(1+\alpha)\mathcal{L}_{\text{rtn}}(\pi_{\text{ref}},\mathcal{D}_{\text{rtn}})\quad\alpha>0.$ (3)
which ensures that the updated model $\pi$ does not degrade retention performance by more than a factor of $\alpha$ relative to the reference model $\pi_{\mathrm{ref}}$.
Unlike scalarization, the constrained formulation explicitly separates the forgetting objective from the retention requirement. This makes the trade-off transparent and easier to interpret: instead of tuning $\lambda$ to balance two competing objectives, the user directly specifies a retention budget $\varepsilon$, and the algorithm maximizes forgetting subject to this constraint.
3.1 Lagrangian Relaxation and the Dual Problem
For the constrained problem  (2), we define the the Lagrangian function:
$\mathcal{L}(\pi,\lambda)=\mathcal{L}_{\text{fgt}}(\pi,\mathcal{D}_{\text{fgt}})+\lambda\left(\mathcal{L}_{\text{rtn}}(\pi,\mathcal{D}_{\text{rtn}})-\varepsilon\right),$
where the Lagrangian multiplier $\lambda\geq 0$ relaxes the hard constraint by a soft penalty. Consider the primal and dual formulations:
[…]
The Primal problem is equivalent to the constrained problem (2), but this primal form is not useful for algorithmic purposes, as the inner maximization over $\lambda$ is unbounded for any fixed $\pi$ that violates the constraint. This motivates the Dual formulation, which, by weak duality [3], finds the largest lower bound on the optimal forgetting loss subject to the constraint. If strong duality holds, the optimal values of both problems coincide, and solving the dual problem yields a solution to the original constrained problem (2).
To ensure zero duality gap, we make two assumptions: (1) convexity: The loss functionals $\pi\mapsto\mathcal{L}_{\text{fgt}}(\pi,\mathcal{D}_{\text{fgt}})$ and $\pi\mapsto\mathcal{L}_{\text{rtn}}(\pi,\mathcal{D}_{\text{rtn}})$ are convex, lower semi-continuous, and defined over the convex policy class $\Pi$; (2) strict feasibility: The constraint is strictly feasible; i.e., there exists $\hat{\pi}\in\Pi$ such that $\mathcal{L}_{\text{rtn}}(\hat{\pi},\mathcal{D}_{\text{rtn}})<\varepsilon$. Under these assumptions, strong duality holds by classical results in convex analysis [32]. This principle underlies a range of recent constrained learning frameworks, including safe reinforcement learning [28], continual learning [10], and constraint-aware LLM fine-tuning via DPO [16].
In our setting, strict feasibility is guaranteed by construction. Specifically, the reference model $\pi_{\text{ref}}$ satisfies the constraint strictly as long as the tolerance parameter $\alpha$ is positive in (3). Hence, strong duality holds as long as the forget and retention losses are convex in the policy $\pi$.
Finite-dimensional parameterization
In practice, the model $\pi$ is parameterized by a finite dimensional parameter $\theta\in\mathbb{R}^{p}$, giving rise to the parameterized dual objective:
$\max_{\lambda\geq 0}\;\min_{\theta\in\Theta}\;\mathcal{L}_{\text{fgt}}({\pi_{\theta}},\mathcal{D}_{\text{fgt}})+\lambda\left(\mathcal{L}_{\text{rtn}}({\pi_{\theta}},\mathcal{D}_{\text{rtn}})-\varepsilon\right).$ (4)
Here, the search space is restricted to $\Pi_{\theta}=\{\pi_{\theta}\mid\theta\in\Theta\}\subseteq\Pi$. While strong duality may not strictly hold in this finite-dimensional, nonconvex setting, modern models are typically sufficiently expressive to approximate the infinite-dimensional problem well [10].
3.2 Proposed Method: Primal-dual with Warm Start
A principled method to solve the above dual problem is dual ascent, which alternates between minimizing the Lagrangian $\mathcal{L}(\theta,\lambda)$ with respect to $\theta$ and applying one step of gradient ascent in $\lambda$ to penalize constraint violation:
$\theta^{+}=\arg\min_{\theta}\mathcal{L}({\pi_{\theta}},\lambda),\quad\lambda^{+}=[\lambda+\eta_{\lambda}(\mathcal{L}_{\text{rtn}}(\pi_{\theta^{+}},\mathcal{D}_{\text{rtn}})-\varepsilon)]_{+},\ \eta_{\lambda}>0.$
The primal update corresponds to minimizing a scalarized objective, while the dual update can be interpreted as dynamically adjusting the trade-off according to the violation $\mathcal{L}_{\text{rtn}}(\pi_{\theta^{+}},\mathcal{D}_{\text{rtn}})-\varepsilon$.
While dual ascent offers strong theoretical guarantees, it typically involves a costly inner-loop optimization to fully minimize the Lagrangian at each step. We propose an efficient variant that performs a single warm-started dual ascent step, followed by lightweight primal-dual updates. The initial iteration fully minimizes $\mathcal{L}({\pi_{\theta}},\lambda_{0})$ with respect to $\theta$. Subsequent iterations alternate between one gradient descent step on $\theta$ and one dual ascent step on $\lambda$, reducing computation through single-step updates while retaining the advantages of dual ascent initialization. This method is detailed in Algorithm 1.
Algorithm 1 Primal-Dual Solver with Warm Starting (Problem (2)
1:Input: Forget set $\mathcal{D}_{\text{fgt}}$, retain set $\mathcal{D}_{\text{rtn}}$, batch sampling algorithm $\mathcal{R}$, reference parameters $\theta_{\mathrm{ref}}$, constraint threshold $\varepsilon$, learning rates $\eta_{\theta},\eta_{\lambda}>0$, initial dual variable $\lambda_{0}\geq 0$, number of warm-up epochs $T_{\mathrm{w}}$, number of primal-dual epochs $T_{\mathrm{pd}}$
2:Output: Primal parameters $\theta^{*}$, dual variable $\lambda^{*}$
3:Initialize: $\theta\leftarrow\theta_{\mathrm{ref}}$, $\lambda\leftarrow\lambda_{0}$
4:for $t=1,\dots,T_{\mathrm{w}}+T_{\mathrm{pd}}$ do
5:     for $d_{\mathrm{fgt}},d_{\mathrm{rtn}}$ in $\mathcal{R}(\mathcal{D}_{\text{fgt}},\mathcal{D}_{\text{rtn}})$ do
6:         $\ell_{\mathrm{f}}(\theta)\leftarrow\mathbb{E}_{(x,y)\in d_{\mathrm{fgt}}}\ell_{\mathrm{fgt}}(\pi_{\theta}(x),y),\quad\ell_{\mathrm{r}}(\theta)\leftarrow\mathbb{E}_{(x,y)\in d_{\mathrm{rtn}}}\ell_{\mathrm{rtn}}(\pi_{\theta}(x),y)$
7:         $\mathcal{L}(\theta,\lambda)\leftarrow\ell_{\mathrm{f}}(\theta)+\lambda(\ell_{\mathrm{r}}(\theta)-\varepsilon)$
8:         $\theta\leftarrow\theta-\eta_{\theta}\nabla_{\theta}\mathcal{L}(\theta,\lambda)$triangleright $\nabla_{\theta}\mathcal{L}(\theta,\lambda)=\nabla_{\theta}\ell_{\mathrm{f}}(\theta)+\lambda\nabla_{\theta}\ell_{\mathrm{r}}(\theta)$
9:         if $t>T_{\mathrm{w}}$then $\triangleright$ Warm-start; Solve the primal problem for a fixed $\lambda$ until epoch $T_{\mathrm{w}}$
10:              $\lambda\leftarrow\left[\lambda+\eta_{\lambda}\left(\ell_{\mathrm{r}}(\theta)-\varepsilon\right)\right]_{+}$triangleright Dual update and project onto $\mathbb{R}_{\geq 0}$
11:         end if
12:     end for
13:end for
14:Return: $\theta$, $\lambda$
While our framework is compatible with a broad class of loss functions proposed in prior unlearning literature, we will focus on specific instantiations of $\mathcal{L}_{\text{fgt}}$ and $\mathcal{L}_{\text{rtn}}$, deferring extensions to alternative losses to the Supplementary Material.
3.3 Retention Loss
For the retain loss $\mathcal{L}_{\text{rtn}}$, we follow established practice and adopt the standard cross-entropy loss:
$\mathcal{L}_{\text{rtn}}(\pi,\mathcal{D}_{\text{rtn}})=\mathbb{E}_{(x,y)\in\mathcal{D}_{\text{rtn}}}\left[CE(\pi^{\mathrm{logits}}(y|x),y)\right]=\mathbb{E}_{(x,y)\in\mathcal{D}_{\text{rtn}}}\left[-\log\left(\pi(y|x)\right)\right],$ (5)
where for a response $y$, the autoregressive model defines the conditional probability as $\pi(y|x)=\prod_{i=1}^{|y|}\pi(y_{i}|x,y_{<i})$, with $\pi(y_{i}|x,y_{<i})$ denoting the likelihood of generating token $y_{i}$ given the input $x$ and the previously generated tokens $y_{<i}$.
3.4 Logit Flattening for Efficient Forgetting
A common heuristic for defining the forget loss $\mathcal{L}_{\text{fgt}}$ is the negative cross-entropy (CE) loss on the forget dataset:
$\mathcal{L}_{\text{fgt}}(\pi,\mathcal{D}_{\text{fgt}})=-\mathcal{L}_{\text{rtn}}(\pi,\mathcal{D}_{\text{fgt}}).$ (6)
However, the CE loss is unbounded above, and directly maximizing it during unlearning often leads to gradient explosion and catastrophic collapse. Notably, CE minimization during pretraining serves as an upper bound surrogate for the 0-1 classification loss. Reversing this objective, by maximizing CE, invalidates this surrogate relationship
To induce high uncertainty in model outputs while avoiding these issues, a more stable alternative is to maximize the entropy of the predictive distribution:
$\mathcal{L}_{\text{fgt}}(\pi,\mathcal{D}_{\text{fgt}})=\mathbb{E}_{(x,y)\in\mathcal{D}_{\text{fgt}}}\left[CE\left(\pi^{\mathrm{logits}}(y\mid x),\tfrac{1}{V}\mathbf{1}\right)\right],$
where $\mathbf{1}\in\mathbb{R}^{V}$ is the all-ones vector and $V=|\mathcal{Y}|$ is the vocabulary size. This loss encourages predictions close to the uniform distribution, and can be viewed as an entropy maximization strategy that suppresses memorized responses by flattening the output distribution.
While effective, entropy-based losses suffer from vanishing gradients as the output distribution approaches uniformity, which can slow down convergence and cause instability in later stages of training. Moreover, they require the log-softmax computation, which can be numerically sensitive and costly in large vocabulary settings.
Logit-margin flattening.
We propose an alternative objective that directly penalizes peakedness in the model’s pre-softmax logits. Given logits $\pi^{\mathrm{logits}}(y_{t}|x,y_{<t})$ to input pair $(x,y)\in\mathcal{D}_{\text{fgt}}$, the proposed logit-margin flattening loss is:
$\mathcal{L}_{\text{fgt}}^{\mathrm{LM}}({\pi_{\theta}},\mathcal{D}_{\text{fgt}}):=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{fgt}}}\left[\frac{1}{|y|}\sum_{t=1}^{|y|}\left(\max_{k}\pi^{\mathrm{logits}}(y_{t}|x,y_{<t})_{k}-\frac{1}{V}\sum_{k=1}^{V}\pi^{\mathrm{logits}}(y_{t}|x,y_{<t})_{k}\right)^{2}\right].$
Minimizing this loss drives the logit vector toward a constant (i.e., uniform after softmax), effectively flattening the predictive distribution. Zero loss is achieved if and only if all logits are equal, implying maximal entropy without computing it explicitly. This logit flattening loss offers several benefits over traditional entropy maximization: (1) It avoids log-softmax operations and relies only on max and mean computations over logits, improving numerical stability and reducing runtime overhead in large vocabulary models; (2) The loss maintains nonzero gradients even when predictions are near uniform, enabling more efficient convergence; (3) The loss is convex in the logits $z$, and therefore compatible with convex surrogate models or linear classifiers. This preserves the strong duality properties required by our constrained optimization framework; and (4) The logit margin directly bounds the model’s maximum softmax probability. This is established next.
Proposition 3.1.
If the logit margin satisfies
$\max_{k}\pi^{\mathrm{logits}}(y_{t}|x,y_{<t})_{k}-\frac{1}{V}\sum_{k=1}^{V}\pi^{\mathrm{logits}}(y_{t}|x,y_{<t})_{k}\leq\delta,$
then the maximum softmax probability is upper-bounded as
$\max_{k}\pi(y_{t}|x,y_{<t})_{k}\leq\left(1+(V-1)\,\exp(-\frac{V}{V-1}\,\delta)\right)^{-1}=\frac{1}{V}(1+\delta)+O(\delta^{2})$
Moreover, a key advantage of our approach is its explicit control over the model’s output distribution on $\mathcal{D}_{\text{fgt}}$, unlike prior methods such as NPO [45], SimNPO [11], and Gradient Ascent [41], which lack such guarantees. This control contributes to the stability of our method by anchoring it to a well-defined target distribution, a benefit also noted in prior work on stable unlearning [7, 18].
4 Experiments
Datasets:
We evaluated our unlearning methodology on two established benchmarks: TOFU and MUSE [25, 33, 8]. The TOFU dataset consists of 200 diverse fictional author profiles, each containing 20 question-answer pairs. A designated subset of these profiles, known as the forget set, serves as the target for unlearning. In the main experiments, we choose to forget the subset Forget10 and defer Forget05 and Forget01 to the Supplementary Material. The MUSE benchmark focuses on unlearning in two real-world contexts: Books and News. Due to space constraints, we consider the News subset in the main experiments and defer Books to the Supplementary Material. The News subset is a collection of BBC news articles collected after August 2023.
Models:
To establish the applicability of the methods, we test our methods across a wide scale of models. We include LLAMA 2 7B, LLAMA 2 13b [37], LLAMA 3.1 8B, LLAMA 3.2 1B, LLAMA 3.2 3B [13], and Gemma 7B [36]. We utilize pretrained instruct versions of these models whenever available [^1] We utilize pretrained and finetuned models through HuggingFace. The models are then finetuned on the desired sets to provide our starting checkpoints. See the Supplementary Material for more information.
Methods:
We compare our method, Primal-Dual Unlearning (PDU), against several baselines. The first is the target model that has been trained on $\mathcal{D}_{\text{rtn}}\cup\mathcal{D}_{\text{fgt}}$. Second, we consider an ideal model that has only been trained on $\mathcal{D}_{\text{rtn}}$. Next, we turn to several established methods: GradDiff [41], DPO [25], NPO [45], SimNPO [11], and RMU [21]. For all the baselines, we use $\lambda=1$ as the trade-off parameter.
Evaluation:
To evaluate the effectiveness of the methods, we utilize several established metrics and calculate harmonic means of them to yield single statistics. More specifically, for the TOFU dataset, we utilize model utility and forget success.
  • Model Utility: Established in [25], model utility is a harmonic mean of several likelihood and ROUGE scores [23] calculated over $\mathcal{D}_{\text{rtn}}$ and other holdout sets.
  • Forget Success: We define this metric as the harmonic mean over $1-$ the likelihood on $\mathcal{D}_{\text{fgt}}$, $1-$ the ROUGE score on $\mathcal{D}_{\text{fgt}}$, and the truth ratio on $\mathcal{D}_{\text{fgt}}$. For metric definitions see [25].
For the MUSE dataset, we utilize the metrics retain ROUGE and forget ROUGE.
  • Retain ROUGE: From [33] (KnowMem($\pi$, $\mathcal{D}_{\text{rtn}}$)), the ROUGE score over knowledge on $\mathcal{D}_{\text{rtn}}$.
  • Forget ROUGE: The harmonic mean of KnowMem($\pi,\mathcal{D}_{\text{fgt}}$) and VerbMem($\pi,\mathcal{D}_{\text{fgt}}$) defined in [33].
In addition to the aforementioned traditional automatic metrics, we employ an LLM-based evaluation framework to assess the success of unlearning and knowledge retention. This method leverages an LLM acting as a structured judge to evaluate generated responses. We task the LLM with judging generated texts with respect to a ground truth response on several avenues and prompt the judge to score each metric from 0 to 10:
  • For forgetting tasks: Knowledge Removal, Verbatim Removal, Fluency.
  • For retention tasks: Retention Score, Accuracy, Relevance, Fluency.
We summarize these results into four metrics: forget score, retain score, fluency, and relevance, where scoring higher is better on all metrics. The details of the metrics and the prompt input to the judge can be found in the Supplementary Material.
Results:
The results of our experiments are reported in Table 1 and Table 2. For unlearning, it is critical to take into account the involved metrics at the same time and to not judge based on single metrics. That is, a model can have a perfect forgetting score but at the cost of catastrophic forgetting severely hampering the models utility.
We can see in Table 1 that our method consistently outperforms all other methods across various scales and models by achieving the highest forget successes whilst retaining the best or second best model utilities. We see similar exceptional performance from our methodology on the LLM judged metrics, except for the Fluency metric. Upon further examination, it becomes clear that this is an artifact of the success of the unlearning algorithm. That is, on the forget set the model’s knowledge has been purged and the model abstains from making coherent predictions. Importantly, it should be noted that the model fluency on the other tasks is unaffected. Due to space constraints, we defer the detailed LLM judge statistics to the Supplementary Material.
We see that for the larger 7B and 8B models, GradDiff has a forget success of 00 but an LLM judged forget score of $10$. Studying the generations of the models, we see that the models unlearned via GradDiff abstain from producing any text when prompted with prompts from $\mathcal{D}_{\text{fgt}}$. As such the truth ratio on $\mathcal{D}_{\text{fgt}}$ is essentially 00 and yields a 00 harmonic mean for the forget success. Due to this behaviour, the judge gives a complete forgetting score to the model. However, unlike our method, we see that GradDiff suffers from this artifact in its utility and also the other LLM judged metrics.
We further see that our method performs competitively on the more complex task of MUSE-News per Table 2. That is, our method provides a viable pareto point that provides unique retain and forget ROUGE scores.
Table 1 and Table 2 further point to an important observation: the traditional metrics used for assessing task success, i.e., metrics such as model utility and forget success, are generally indicative of real success, as outlined by the correlation that we see with the LLM judged metrics. Without the LLM judged metrics, it was not clear if metrics such as the likelihood of generating the prompt-response pair $(x,y)\in\mathcal{D}_{\text{fgt}}$ or the ROUGE score would be real indicators of the successful unlearning. The LLM judged metrics show that this is generally the case and classical metrics are still useful indicators of a model’s capabilities.
Table 1: Performance on the TOFU dataset (forget10/retain90) with different unlearning methods and models. Model utility and forget success are bounded in $[0,1]$ whereas the LLM Judged metrics are in $[0,10]$. For all metrics, larger numbers are better. We bolden the best results and underline the runner-ups.
Method Model Forget LLM Judged
Utility Success Forget Score Retain Score Fluency Relevance
Llama 3.2 1B target 0.595 0.194 1.643 8.235 9.695 9.405
retrained 0.590 0.691 7.569 8.464 9.676 9.428
GradDiff 0.434 0.616 7.001 5.748 8.413 8.277
DPO 0.561 0.603 9.231 7.390 9.349 8.678
NPO 0.475 0.672 6.695 5.686 9.012 8.643
SimNPO 0.596 0.248 2.659 8.250 9.646 9.368
RMU 0.570 0.689 7.973 7.415 8.410 9.003
PDU (Ours) 0.602 0.740 8.556 7.885 7.988 9.209
Llama 3.2 3B target 0.660 0.083 0.593 9.159 9.830 9.732
retrained 0.645 0.694 7.673 9.101 9.734 9.731
GradDiff 0.529 0.583 6.766 6.546 8.196 8.470
DPO 0.609 0.540 8.630 8.292 9.415 9.023
NPO 0.514 0.676 6.880 7.184 9.306 8.825
SimNPO 0.653 0.196 1.839 8.898 9.751 9.657
RMU 0.644 0.561 5.966 8.348 9.502 9.469
PDU (Ours) 0.680 0.914 9.558 8.809 7.760 9.617
Llama 3.1 8B target 0.628 0.013 0.0926 9.642 9.904 9.894
retrained 0.649 0.693 7.505 9.646 9.794 9.874
GradDiff 0.626 0 10 8.247 7.257 9.169
DPO 0.497 0.596 9.501 5.345 9.020 6.160
NPO 0.652 0.739 8.329 8.588 9.360 9.509
SimNPO 0.603 0.481 4.630 8.983 9.691 9.698
RMU 0.657 0.900 9.925 9.626 7.969 9.867
PDU (Ours) 0.725 0.960 9.985 9.277 7.717 9.793
Gemma 7B target 0.638 0.0342 0.305 8.655 9.818 9.558
retrained 0.642 0.670 7.623 8.551 9.665 9.552
GradDiff 0.461 0 9.988 4.720 6.766 7.458
DPO 0.488 0.591 7.760 6.728 9.283 8.772
NPO 0.543 0.744 8.631 7.027 9.363 8.873
SimNPO 0.547 0.493 5.901 7.226 9.496 8.963
RMU 0.633 0.630 9.785 8.351 7.656 9.453
PDU (Ours) 0.602 0.933 9.996 7.323 7.303 9.023
Table 2: Performance on the MUSE-News dataset with different unlearning methods and two large scale models. ROUGE scores are bounded in $[0,1]$ whereas the LLM Judged metrics are in $[0,10]$. We bolden the best results and underline the runner-ups.
Method Retain $\uparrow$ Forget $\downarrow$ LLM Judged $\uparrow$
ROUGE ROUGE Forget Success Retain Score Fluency Relevance
Llama 2 7B target 0.555 0.610 2.428 5.810 9.083 8.760
retrained 0.560 0.250 6.905 5.460 9.030 8.670
GradDiff 0.482 0.331 4.300 5.355 8.783 8.500
NPO 0.455 0.318 4.688 4.545 8.687 7.930
SimNPO 0.516 0.573 2.748 5.490 9.033 8.550
RMU 0.460 0.418 4.398 4.855 8.887 8.060
PDU (Ours) 0.397 0.290 5.550 4.040 7.767 7.630
Llama 2 13B target 0.430 0.632 2.695 5.075 9.193 8.31
retrained 0.395 0.255 6.948 4.440 8.920 7.780
GradDiff 0.488 0.287 5.648 5.410 8.777 8.360
NPO 0.420 0.403 4.315 5.015 9.080 8.340
SimNPO 0.448 0.440 4.153 5.375 8.877 8.320
RMU 0.232 0.194 7.865 3.025 8.173 6.640
PDU (Ours) 0.452 0.289 5.050 4.795 8.427 8.050
5 Conclusion
We presented a principled framework for unlearning in Large Language Models by casting the problem as a constrained optimization task. This formulation separates the forgetting and retention objectives, providing explicit control over each. To enable stable and efficient forgetting, we introduced a logit-margin flattening loss that avoids the pitfalls of entropy maximization while encouraging uniform predictive distributions on the forget set. Our scalable primal-–dual solver enforces the retention constraint and exposes the forgetting-–utility trade–off through interpretable dual dynamics. Empirical evaluations on TOFU and MUSE benchmarks demonstrate that our method effectively suppresses memorized responses while preserving retained capabilities, often matching full retraining at a fraction of the cost.
Our work opens several directions for future investigation. First, due to the resource-intensive nature of LLMs, we were unable to conduct extensive hyperparameter tuning; it is possible that further gains could be achieved with careful calibration. Second, we observed a slight reduction in generation fluency on the easier TOFU task under our method, potentially attributable to the strong uniformity induced by logit flattening. Addressing this through regularization or hybrid losses is an interesting direction. Third, while our method is designed to remove specific information, we do not study the resilience of the resulting model to relearning attacks or jailbreak attempts. Finally, a compelling extension would assess the robustness of our framework under continual unlearning requests, where new forget sets arrive over time.